Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

East Antrim Rifle Club Festival 2001

Distribution

Dr. Trevor West DUCAC Chairman
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue East Antrim Rifle Club
Date 7-9 October, 2001
Discipline 50m Prone Rifle, 10m Air Rifle, 25m smallbore pistol

DURC Attendees

John Keeney Ray Reilly

Alan Smyth

Scores 10m Air Rifle

John Keeney 477
Alan Smyth 485

10m Air Pistol

John Keeney 426
Alan Smyth 477

50m Prone Rifle

John Keeney 543
Ray Reilly 575
Alan Smyth 540

25m Smallbore Pistol

John Keeney 303
Alan Smyth 427

Report

Following the successful Freshers’ Week membership drive, three members of DURC attended the EARC Festival. Due to complications with paperwork, a change of dates and the level of work during Freshers’ Week, our original planned team of eight was seriously whittled down.

John and Alan shot very disappointing scores due to extreme lighting conditions early saturday morning, where the early morning sun was directly behind the target thereby eliminating almost all contrast in the sight picture.

It’s hoped that next year we will be able to send more people to EARC for the competition, this year saw a lot of factors come together to prevent this.

Misé le meas,

Mark Dennehy, DURC Secretary